OpenStudy (anonymous):

Simplify the radical expression. -3 the square root of 180h^4 (i will draw it out)

OpenStudy (anonymous):



OpenStudy (anonymous):

\[{-3}\sqrt{180h^4}\]

OpenStudy (campbell_st):

\[-3\sqrt{180h^4}= -3\times \sqrt{36\times5\times(h^2)^2}\] or \[-3\times \sqrt{36}\times \sqrt{(h^2)^2}\times \sqrt{5}\] \[-3\times6\times h^2 \times \sqrt{5}\] answer \[-18h^2\sqrt{5}\]

OpenStudy (anonymous):

factorize 180 = 2*2*3*3*5 so \[{-3}\sqrt{2^2\times3^2\times5\times{h^4}}={-3}\times2\times3\times{h^2}\sqrt{5}=-18h^2\sqrt{5}\]